HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

QUANTUM CHEMICAL AND MICROWAVE SPECTROSCOPIC ANALYSIS OF 1-CYANOPROPENE AND 1-CHLOROPROPENE

\underline{T. A. NGUYEN^{1}}, I. KLEINER^{1}, M. SCHWELL^{2}, H. V. L. NGUYEN^{2,3}, 1Université Paris Cité and Univ Paris Est Créteil, CNRS, LISA, 75013 Paris, France,^{2}Univ Paris Est Créteil and Université Paris Cité, CNRS, LISA, 94010 Créteil, France,^{3}Institut Universitaire de France (IUF), 75231 Paris, France

The 1-cyanopropene and 1-chloropropene molecules are methylated derivatives of propene which contain a nitrogen and a chlorine atom, respectively. They belong to a group of important substances playing a major role in astrophysics^{1} and atmospheric chemistry. The aim of this work is to study their gas phase structures using a combination of pulsed molecular jet Fourier transform microwave spectroscopy and quantum chemical calculations. The microwave spectra were measured in the frequency range from 2 to 40 GHz. Quantum chemical calculations were performed at the B3LYP-D3BJ/6-311++G(d,p), MP2/6-311++G(d,p), and MP2/6-31G(d,p) levels of theory to obtain the optimized molecular geometries. Due to internal rotation of the methyl group, all rotational transitions split into an A and an E torsional species. In addition, each of these two torsional species contains further hyperfine splittings arising from the quadrupole coupling of the nitrogen as well as the chlorine nucleus. The complex spectral splittings were analyzed and modeled using the XIAM^{2} and the BELGI^{3}^{,}^{4} programs to obtain experimental molecular parameters and validate the quantum chemical results.

1doi:10.1051/0004-6361/202244255, J. Cernicharo, R. Fuentetaja, C. Cabezas, M. Agundez, N.

Marcelino, B. Tercero, J. R. Pardo, and P. de Vicente, Discovery of Five Cyano Derivatives of Propene with the QUIJOTE Line Survey, A&A. 663, L5 (2002).

2doi:10.1515/zna-1996-0807, H. Hartwig and H. Dreizler, The Microwave Spectrum of trans-2,3-Dimethyloxirane in Torsional Excited States, Z. Naturforsch. 51a, 923 (1996).

3doi:10.1006/jmsp.1994.1047, J.T. Hougen, I. Kleiner, and M. Godefroid, Selection Rules and Intensity Calculations for a C_{s} Asymmetric Top Molecule Containing a Methyl Group Internal Rotor, J.

Mol. Spectrosc. 163, 559 (1994).

410.1021/acs.jpca.6b02111, R. Kannengießer, W. Stahl, H.V.L. Nguyen, and I. Kleiner, ^{14}N Nuclear Quadrupole Coupling and Methyl Internal Rotation in N-tert-Butylacetamide As Observed by Microwave Spectroscopy, J. Phys. Chem. A. 120, 3992 (2016).

p-number: p172  Submitted on Thu, 15 Jun 2023 12:10:07 +0200